OpenStudy (anonymous):

please help me.....Mr. Hemley stands next to Illinois Centennial Monument at Logan Square in Chicago and casts a shadow that is 18 feet long. The shadow of the moniment is 204 feet long. If Mr. Hemley is 6 feet tall, how tall is the monument?

OpenStudy (anonymous):

think is is 68.67 take 206 * 6=1236 then / that by 18

OpenStudy (anonymous):

then round it

OpenStudy (anonymous):

sorry...but I could not get you...please explain... :(

OpenStudy (anonymous):

the answer is 68.67

OpenStudy (anonymous):

but how could you get the answer?? ...I don't have a clue...please explain the process

OpenStudy (anonymous):

68 is answer

OpenStudy (anonymous):

The shadow of the moniment is 204 feet X Mr. Hemley is 6 feet tall, divide Logan Square in Chicago and casts a shadow that is 18 feet long the the answer you get by multiplying 204X6=1224/18=68

OpenStudy (anonymous):

Way i remember doing it in math

OpenStudy (anonymous):

ohh....thanks a lot... :)

OpenStudy (anonymous):

was that the answer hopefuly could help you out

OpenStudy (anonymous):

yeah...it helped me... :)

OpenStudy (anonymous):

good glad to help